one bacterial culture is 18 days older than a second culture. In 1o days the first culture will be twice the age of the second. what are the present ages of the two cultures??
Answer by stanbon(48545)

The problem tells you how the two cultures ages are related NOW:
1st culture is x+18 days; 2nd culture is x days
It also tells you to consider their ages 10 days from NOW:
1st culture will be x+18+10; 2nd culture will be x+10
Then it tells you the 1st culture is twice as old as the 2nd 10 days
from NOW. This gives you the equation you need.
Equation: 1st = 2 (2nd)
x+28 = 2(x+10)
Solve for "x" and "x+8"
Cheers,
Stan H.